# DMIS: Dynamic Mesh-based Importance Sampling for Training Physics-Informed Neural Networks

### 17 Jan 2024 v2

@@ Introduction
Datasets for PDE residuals, initial conditions and boundary conditions are denoted by $N_{f}$, $N_{i}$ and $N_{b}$, respectively. In general, mini-bitches are uniformly sampled from $N_{f}$, $N_{i}$, and $N_{b}$, respectively. According to the Monte Carlo approximation, we can introduce a more efficient sampling method. Because boundary conditions and initial conditions are penalty terms, we only introduce importance sampling into the sampling of $N_{f}$. The loss of PDE residuals $\mathcal{L}_{f}$ combined with importance sampling is shown as:
@@ Appendix A Appendix A: Supporting Theory
where $\alpha_{\mathrm{x}}$ is a sample weight, $\bm{\theta}^{(t)}$ and $\bm{\theta}^{(t+1)}$ are parameters at iterations $t$ and $t+1$, respectively, and $\eta$ is learning rate. For uniform sampling, $\alpha_{\mathrm{x}}=1$. Assuming that the sampling probability of $\mathrm{x}$ is $q_{\mathrm{x}}$, with the condition of unbiased estimation, the sample weight $\alpha_{\mathrm{x}}=1/(|N|q_{\mathrm{x}})$. We aim to find the best sampling probability $q_{\mathrm{x}}$ to make the training convergence rate fastest.
@@
The residual of boundary conditions is an integral over the boundary of domain:
@@
Equation 22 can be approxmated as a loss function of data points:
@@
where $\mathcal{L}$ is the total loss, $\mathcal{L}_{f}$, $\mathcal{L}_{i}$, and $\mathcal{L}_{b}$ are the losses of PDE residuals, initial conditions, and boundary conditions, respectively. The datasets of these losses are denoted by $N_{f}$, $N_{i}$ and $N_{b}$, respectively.
@@ Appendix B Appendix C: Additional Experiment
Table 8 summarizes hyper-parameters for each additional benchmark. The Adam optimizer is used on all additional benchmarks. For PINN-N, we uniformly sample 10,000 seeds within the domain, and other parameters are consistent with Table 8. Experiments on each additional benchmark have same random seed.
@@
Figure 7(a, b, c) report the prediction error by PINN-O, PINN-N, and PINN-DMIS for solving the diffusion equation. Compared with other state-of-the-art methods, DMIS has best performance. The snapshots in Figure 7 show that DMIS has the best prediction results in the whole domain. Table 9 summarizes the evaluation results of prediction accuracy for solving Diffusion Equation. DMIS has the best performance on all evaluation metrics. Figure 7(d) reports the convergence curves for solving the Diffusion Equation. DMIS converges fastest and has the lowest training error. However, the loss value fluctuates greatly in DMIS. We speculate that the loss value fluctuates greatly due to the unstable sample weights. Affected by the complex initial conditions, the triangular mesh updates frequently during training. Table 10 compares the convergence behaviors. DMIS has the fastest convergence speed.
@@
Figure 8(a, b, c) report the prediction errors for solving the Allen-Cahn Equation. DMIS has the lowest prediction error. Figure 8(e, f, g) report the snapshots of prediction at t=0.8347s. Compared with PINN-O and PINN-N, DMIS has better prediction results on the position with large gradient. Table 9 summarizes the evaluation results for solving the Allen-Cahn Equation. DMIS has the best performance on all evaluation metrics. The convergence curves for solving the Allen-Cahn Equation are reported in Figure 8(d). DMIS does not show obvious advantages. The result shows that DMWE fails to generate an effective mesh. The failure reason are analyzed in detail in next section. Table 11 reports the convergence behaviors for solving the Allen-Cahn Equation. These three methods have similar convergence behavior and DMIS has a slender lead convergence speed.
Figure 7: The Diffusion Equation. (a, b, c) The prediction errors of PINN-O, PINN-N and PINN-DMIS, respectively. (d) Convergence curves. (e, f, g) show the prediction (red) and ground-truth (blue).
@@ Appendix C Appendix D: Limitations of DMIS
Figure 8: The Allen-Cahn Equation. (a, b, c) The prediction errors of PINN-O, PINN-N and PINN-DMIS, respectively. (d) Convergence curves. (e, f, g) show the prediction (red) and ground-truth (blue).
@@
In addition to the limitation brought by importance sampling, the triangular mesh also affects the effect of DMIS. For DMIS, the instability of the loss value distribution leads to frequent updating of the triangular mesh, which affects the training efficiency and the stability of training. In solving the diffusion equation, the mesh updates frequently, and the loss value fluctuates greatly, as shown in Figure 7(d). These defects can be alleviated by setting an appropriate updating threshold. The selection of the threshold mainly depends on experience. Fortunately, our experiments show that the impact of threshold is relatively stable within a reasonable range $[0.3,0.6]$ and we recommend setting the threshold $\gamma$ to 0.4.
In this paper, we only verify the excellent performance and generalization ability of DMIS on five widely-used benchmarks. We will verify the applicability of DMIS on more complex PDEs, for example high-dimensional PDEs, in the future.

### 18 Jan 2024 v3

@@ Introduction
Datasets for PDE residuals, initial conditions and boundary conditions are denoted by $N_{f}$, $N_{i}$ and $N_{b}$, respectively. In general, mini-batches are uniformly sampled from $N_{f}$, $N_{i}$, and $N_{b}$, respectively. According to the Monte Carlo approximation, we can introduce a more efficient sampling method. Because boundary conditions and initial conditions are penalty terms, we only introduce importance sampling into the sampling of $N_{f}$. The loss of PDE residuals $\mathcal{L}_{f}$ combined with importance sampling is shown as:
@@ Appendix A Appendix A: Supporting Theory
where $\alpha_{\mathrm{x}}$ is a sample weight, $\bm{\theta}^{(t)}$ and $\bm{\theta}^{(t+1)}$ are parameters at iterations $t$ and $t+1$, respectively, and $\eta$ is learning rate. For uniform sampling, $\alpha_{\mathrm{x}}=1$. Assuming that the sampling probability of $\mathrm{x}$ is $q_{\mathrm{x}}$, with the condition of unbiased estimation, the sample weight $\alpha_{\mathrm{x}}=1/(|N|q_{\mathrm{x}})$. We aim to find the best sampling probability $q_{\mathrm{x}}$ to make the training convergence rate the fastest.
@@
The residual of boundary conditions is an integral over the boundary of the domain:
@@
Equation 22 can be approximated as a loss function of data points:
@@
where $\mathcal{L}$ is the total loss, $\mathcal{L}_{f}$, $\mathcal{L}_{i}$, and $\mathcal{L}_{b}$ are the losses of PDE residuals, initial conditions, and boundary conditions, respectively. The datasets of these losses are denoted by $N_{f}$, $N_{i}$, and $N_{b}$, respectively.
@@ Appendix B Appendix C: Additional Experiment
Table 8: The hyper-parameters used for each benchmark. $|S|$ is the set size of mesh points, $\gamma$ is the mesh update threshold, and $\beta$ is the hyper-parameter of reweighting. $|N_{f}|$, $|N_{i}|$, and $|N_{b}|$ are the dataset size of PDE residuals, initial conditions, and boundary conditions, respectively.
@@
Table 8 summarizes hyper-parameters for each additional benchmark. The Adam optimizer is used on all additional benchmarks. For PINN-N, we uniformly sample 10,000 seeds within the domain, and other parameters are consistent with Table 8. Experiments on each additional benchmark have the same random seed.
@@
Figure 7(a, b, c) report the prediction error by PINN-O, PINN-N, and PINN-DMIS for solving the diffusion equation. Compared with other state-of-the-art methods, DMIS has the best performance. The snapshots in Figure 7 show that DMIS has the best prediction results in the whole domain. Table 9 summarizes the evaluation results of prediction accuracy for solving Diffusion Equation. DMIS has the best performance on all evaluation metrics. Figure 7(d) reports the convergence curves for solving the Diffusion Equation. DMIS converges fastest and has the lowest training error. However, the loss value fluctuates greatly in DMIS. We speculate that the loss value fluctuates greatly due to the unstable sample weights. Affected by the complex initial conditions, the triangular mesh updates frequently during training. Table 10 compares the convergence behaviors. DMIS has the fastest convergence speed.
@@
Figure 8(a, b, c) report the prediction errors for solving the Allen-Cahn Equation. DMIS has the lowest prediction error. Figure 8(e, f, g) report the snapshots of prediction at t=0.8347s. Compared with PINN-O and PINN-N, DMIS has better prediction results on the position with a large gradient. Table 9 summarizes the evaluation results for solving the Allen-Cahn Equation. DMIS has the best performance on all evaluation metrics. Figure 8(d) reports the convergence curves for solving the Allen-Cahn Equation. DMIS does not show obvious advantages. The result shows that DMWE fails to generate an effective mesh. The failure reasons are analyzed in detail in the next section. Table 11 reports the convergence behaviors for solving the Allen-Cahn Equation. These three methods have similar convergence behavior and DMIS has a slender lead convergence speed.
Figure 7: The Diffusion Equation. (a, b, c) The prediction errors of PINN-O, PINN-N, and PINN-DMIS, respectively. (d) Convergence curves. (e, f, g) show the prediction (red) and ground truth (blue).
@@ Appendix C Appendix D: Limitations of DMIS
Figure 8: The Allen-Cahn Equation. (a, b, c) The prediction errors of PINN-O, PINN-N, and PINN-DMIS, respectively. (d) Convergence curves. (e, f, g) show the prediction (red) and ground truth (blue).
@@
In addition to the limitation brought by importance sampling, the triangular mesh also affects the effect of DMIS. For DMIS, the instability of the loss value distribution leads to frequent updating of the triangular mesh, which affects the training efficiency and stability of training. In solving the diffusion equation, the mesh updates frequently, and the loss value fluctuates greatly, as shown in Figure 7(d). These defects can be alleviated by setting an appropriate updating threshold. The selection of the threshold mainly depends on experience. Fortunately, our experiments show that the impact of the threshold is relatively stable within a reasonable range $[0.3,0.6]$ and we recommend setting the threshold $\gamma$ to 0.4.
In this paper, we only verify the excellent performance and generalization ability of DMIS on five widely-used benchmarks. We will verify the applicability of DMIS on more complex PDEs, for example, high-dimensional PDEs, in the future.